Strong base is dissolved in 625 mL of 0.400 M weak acid (Ka = 3.76 × 10-5) to make a buffer with a pH of 4.09. Assume that the volume remains constant when the base is added.
Answer #1

first calculate pKa from Ka value using

pKa = -logKa = -log[3.76 × 10-5]   = 4.4

pH = pka + log (base/acid)

4.09 = 4.4 + log (base/acid)

log (base/acid) = 4.09 - 4.4

log (base/acid) = -0.31

(base/acid) = 10-0.31

(base/acid) = 0.500

moles of the acid = Molarity x volume in liters = 0.4 x 0.625L   = 0.25 moles acid

HA + XOH ----> XA + H2O

so here no of moles of salt formed no of moles of strong base

lets base = x

x / 0.25-x   = 0.5

x = 0.125 - 0.5x

1.5x = 0.125

x = 0.125 / 1.5

x = 0.083 M

moles = M x volume in liters = 0.083 x 0.625 L = 0.0518 Moles

i calculated both moles and molarity find out whatever you want